# Level-Set Parameters: Novel Representation for 3D Shape Analysis

### v1

@@ 3.2 Dataset Construction
As introduced in Eq. (1), we decompose the level-set parameters $\boldsymbol{\theta}$ into $\boldsymbol{\mu}+\Delta\boldsymbol{\theta}$. This is inspired by the reparameterization trick in variational autoencoders [70]. It emulates a normal distribution with expectation $\boldsymbol{\mu}$ and homogeneous standard deviation $1$. We follow this decomposition to construct a dataset of level-set parameters with shape transformations in two stages.
@@
with $\{\Delta\theta_{1}^{mn}\}$ being pose-dependenet. $\odot$ denotes the Hadamard product. The tanh function constrains $\Delta\boldsymbol{\theta}$ to be in the range $[-1,1]$. We train the above parameters similarly based on Algorithm 1. The major difference from stage one is that all shapes in a batch become clones of the individual shape being fitted. In addition, note that $\boldsymbol{\theta}=\boldsymbol{\mu}+\Delta\boldsymbol{\theta}$ in the Algorithm in this stage.
@@
Why not meta-learning like MetaSDF? MetaSDF [15] applies the meta-learning technique [68] to per-category shape reconstruction of DeepSDF. It introduces a large number of additional parameters (i.e., per-parameter learning rates) to train the network, which is computation-intensive. Besides, its three gradient updates of the network during inference stage often result in unsatisfactory surface quality [71]. Moreover, the loss function we employ in Eq. (7) for unsupervised surface reconstruction requires input gradients of the SDF network to be computed with backpropagation at every iteration, prohibiting the application of MetaSDF or MAML [68].
@@ 4.1 Encoder-based Semantic Learning
We format the level-set parameters into multiple tensors for shape analysis. Specifically, the weights and biases in the first SDF layer are concatenated into a tensor of size $256{\times}4$. Further, we concatenate all parameters from layer 2 to 7 into a tensor of size $6{\times}256{\times}257$, while zero-padding is applied to the parameters of the skipping layer. Regarding the final SDF layer, its parameters are combined into a tensor of size $1{\times}257$. Thus, each shape surface is continuously represented as a tuple of three distinct tensors $(\Theta_{1}{\in}\mathbb{R}^{256{\times}4},\Theta_{2}{\in}\mathbb{R}^{6{\times}256{\times}257},\Theta_{3}{\in}\mathbb{R}^{1{\times}257})$. See supplemtary for an illustration.
@@
The proposed semantic learning network has three branches, each processing a different component in the input tensors ($\Theta_{1},\Theta_{2},\Theta_{3}$), as depicted in Fig. 2(a). It builds upon the BaseNet and BasePool blocks shown in Fig. 2(b), as well as the PECat block in Fig. 2(c). The BaseNet block comprises two fully connected layers followed by batch normalization [72]. The first layer applies ReLU activation. For batch normalization, we flatten the first two dimensions of the input tensor. For example, the resulting dimensions for $\Theta_{1}$, $\Theta_{2}$, $\Theta_{3}$ will be $1024$, $1536{\times}257$, $257$. The BasePool block involves BaseNet followed by average pooling, which is required in processing the three-dimensional $\Theta_{2}$.
The PECat block differentiate level-set parameters from different SDF layers in tensor $\Theta_{2}$. It applies positional encoding [73] to the layer index $\ell{\in}\{2,3,4,5,6,7\}$, and concatenates the encoded features $\text{PE}(\ell)$ to their corresponding layer parameters in $\Theta_{2}$. Let $d_{pe}$ be the dimension of $\text{PE}(\ell)$. Thus, the tensor $outA$ in Fig. 2(c) has dimension of $6{\times}256{\times}(257{+}d_{pe})$. We process it with BasePool2A. Besides, to extract semantics from the 2nd dimension of $\Theta_{2}$, we transpose $\Theta_{2}$ into $\Theta_{2}^{\intercal}{\in}{\mathbb{R}^{6{\times}257{\times}256}}$ and concatenate it with $\text{PE}(\ell)$ to get a tensor $outB$ of size $6{\times}257{\times}(256{+}d_{pe})$, processed by BasePool2B.
The BaseNet1 and BaseNet3 modules compute branch features $\mathbf{a}_{1}{\in}\mathbb{R}^{256}$ and $\mathbf{a}_{3}{\in}\mathbb{R}^{256}$, respectively, based on $\Theta_{1}$ and $\Theta_{3}$. Meanwhile, BasePool2A and BasePool2B each compute features of size $256$, resulting in the concatenated feature $\mathbf{a}_{2}{\in}\mathbb{R}^{512}$. These features are further concatenated to form a unified global surface feature $\mathbf{s}_{g}{\in}\mathbb{R}^{1024}$, which is applicable to both shape classification and retrieval. We train the network with standard cross-entropy loss. The classifier comprises a single FC layer.
Our analysis considers level-set parameters from all layers of SDF, which forms a complete shape representation. In contrast, prior works utilized partial level-set parameters. For instance, [17] did not include the first layer parameters, while [16] excluded the final layer parameters. Further, given our proposed decomposition of $\boldsymbol{\theta}$ and the fact that $\boldsymbol{\mu}$ is shared by all shapes, we normalize the level-set parameters with $\boldsymbol{\mu}$ and study shape semantics with $\Delta\boldsymbol{\theta}$, the instance parameters of each shape.
@@ 5 Experiment
ShapeNet [23] and Manifold40 [6] provide multi-category 3D shapes that are well-suited for geometric research in computer graphics and robotics. For the interested shape analysis, we construct Level-Set Parameter Data (LSPData) as continuous shape representations for these datasets. The two-stage training of HyperSE3-SDF in § 3.2 is adopted in this process. In the first stage of constructing the LSPData, we utilize 20 shapes from each class in ShapeNet and 7 shapes per class in Manifold40 to train the pose-dependent initialization $\boldsymbol{\mu}$. The point clouds of each shape for SDF are sampled from the shape surfaces in Manifold40 and sourced from previous work [10] in ShapeNet. We did not consider the loudspeaker category in ShapeNet due to the intricate internal structures of the shapes.
@@ 5.1 Understand the Parameter Data
In the proposed dataset construction, we train HyperSE3-SDF for all shapes to create the LSPData with transformations for continuous shape analysis. To validate this approach, we conduct two extra experiments based on the plain SDF network to construct the LSPData, using different settings for $\boldsymbol{\mu}$. In the first experiment, we randomly initialize $\boldsymbol{\mu}$ and apply it to all shapes as [17]. In the second, we learn $\boldsymbol{\mu}$ using the same training samples as those used for HyperSE3-SDF. The training process resembles the algorithm described in § 3.2, except that we fix the pose as $(\mathbf{I},\mathbf{0})$ without sampling.
In this ablation study, we compare the constructed LSPData for five major categories of ShapeNet [7]. We present the t-SNE embeddings [74] of the data constructed by random and learned $\boldsymbol{\mu}$ in Fig. 3. We note that the embeddings for our HyperSE3-SDF in pose $(\mathbf{I},\mathbf{0})$ are similar to Fig. 3(b). It can be seen that the learned $\boldsymbol{\mu}$ significantly enhances shape semantics compared to the randomly initialized counterpart. Table 1 compares the shape classification results of the different LSPData under different rotation setups. We utilize the encoder-based network in § 4.1 for the classification. ‘$(\mathbf{I},\mathbf{0})$’ represents testing with data in reference poses. ‘$z$/SO(3)’ indicates training with data rotated around $z$ axis but testing on data arbitrarily rotated in SO(3) [19]. We obtain the transformed LSPData of plain SDF by applying Eq. (4). Notably, the transformed LSPData from HyperSE3-SDF surpasses those from Euclidean computations, while the normalized LSPData $\Delta\boldsymbol{\theta}$ outperforms the unnormalized $\boldsymbol{\theta}$. We also conduct ablation studies to identify the optimal pose-dependent subset within the level-set parameters for the hypernetwork to generate. Details can be found in the supplementary material.
@@ 5.2 Shape Classification and Retrival
We apply the level-set parameters to semantic analysis on ShapeNet and Manifold40. For time concern, a maximum of 2000 shapes are reconstructed for each class in ShapeNet. We filter the continuous shapes based on their Chamfer distances to ensure surface quality. This results in a dataset of 17,656 shapes for ShapeNet and 10,859 for Manifold40. Figure 4 plots the mean and standard deviation of Chamfer distance (CD1) for each class in ShapeNet, illustrating the surface quality represented by our LSPData. Additionally, the mean and standard deviation of CD1 for all classes in Manifold40 are $0.7102$ and $0.1443$, respectively. Note that our chamfer distances are averaged across multiple random poses due to the introduced surface transformation. Besides, our training set in ShapeNet comprises 200 shapes per class, while in Manifold40, it contains 50% of the shapes per class, up to 200 shapes.
ShapeNet. We conduct pose-related semantic analysis on ShapeNet using rotation setups of $z$/SO(3) and $(\mathbf{I},\mathbf{0})$. Table 2 demonstrates the feasibility of shape analysis based on level-set parameters without dependence on point clouds and meshes [17]. The proposed level-set parameters, LSPData, show comparable performance to point clouds in shape classification of pose $(\mathbf{I},\mathbf{0})$. In the challenging setup of $z$/SO(3), our encoder-based network for LSPData outperforms the rotation-equivariant network VN-DGCNN [18] for point clouds. Conversely, the point cloud-based networks, PointNet [1] and DGCNN [2], yield unsatisfactory results for not addressing rotation equivariance. We note that the equivariance of VN-DGCNN comes with time cost, requiring 9×9$\times$9 × more time for training compared to DGCNN. In contrast, the classification network for LSPData converges rapidly (e.g., within $20$ minutes on the same machine using a single NVIDIA RTX 4090), akin to the observation in [65].
@@
Method pose OA plane bench cab car chair display lamp rifle sofa table phone vessel PointNet [1] $z$/SO(3) 27.34 17.01 33.16 27.41 13.87 29.83 21.88 35.28 61.11 12.36 11.74 22.61 38.13 DGCNN [2] 24.41 81.76 4.82 19.31 43.75 13.60 4.28 59.87 0.00 4.09 4.65 8.38 22.16 VN-DGCNN [18] 90.40 97.01 81.84 93.35 98.35 86.98 83.62 89.05 96.17 89.39 80.95 92.95 93.13 LSPData ($\Delta\boldsymbol{\theta}$) 91.54 97.24 82.36 83.18 98.08 93.43 86.55 89.46 98.15 88.81 85.01 97.21 95.15 PointNet [1] ($\mathbf{I,0}$) 92.48 97.60 86.46 95.02 98.56 94.07 90.59 87.37 98.15 91.81 81.68 91.76 94.18 DGCNN [2] 94.55 99.00 84.35 95.64 99.86 93.94 94.25 93.31 99.15 92.06 88.26 97.87 97.69 LSPData ($\Delta\boldsymbol{\theta}$) 93.46 98.59 84.02 88.16 98.83 94.65 95.11 93.14 99.57 90.56 87.00 94.15 95.75
@@
Manifold40. We also compare the effectiveness of LSPData in shape classification and retrieval with point clouds under exhaustive rotation augmentations, i.e., the SO(3)/SO(3) setup, abbreviated as SO(3) in Table 3. For retrieval, we take features after the average pooling layer to match shapes in the feature space via Euclidean distances. We evaluate the retrieval performance using mean Average Precision (mAP) alongside the top-1/5/10 recalls. It can be noticed that the proposed encoder-based network for LSPData still outperforms the VN-DGCNN [18], DGCNN [2], and PointNet [1] for point clouds. The performance gap between Manifold40 and ShapeNet is mainly attributed to the restricted number of shapes in certain classes of Manifold40.
@@ 5.3 Object Pose Estimation
Data preparation. To prepare the partial point clouds for pose estimation, we select three categories from ShapeNet: airplane, car, and chair, with 10 shapes utilized in each category. For every shape, we create 10 ground-truth transformations with random rotations in the range of $[0,2\pi]$ and translations in the range of $[-0.1,0.1]$. For each transformed shape, we create the partial point cloud from its full point cloud representation, using hidden point removal [75]. This results in 300 pairs that covers partial and full point clouds with limited overlaps and large rotations. We introduce different levels of noise, $\sigma{\in}\{0.01,0.03\}$, and an outlier ratio of 30% to the partial point clouds for further challenges.
We compare the estimation quality of our method, Proposed V1 and V2 (see § 4.2), with optimization-based registration methods, including ICP [54], FGR [55], and TEASER++[56] in Table 4. Their performance is evaluated based on the relative rotation/translation errors (RRE/RTE). See the supplementary for definitions of the two metrics. RRE is reported in degrees, and RTE is scaled by ×$\times 100$. Notably, the proposed method effectively estimates poses with arbitrary rotations from partial-view point clouds, even in the presence of significant noise and outliers. Visualized examples are provided in Fig. 5, with more in the supplementary. We notice that while TEASER++ recovers most poses in the clean data, its performance drops with noise and outliers. ICP and FGR struggle with estimating large rotations even for clean point clouds. If a method fails in simpler settings, we cease testing it on more challenging data. The proposed method takes $\sim$50 seconds to estimate the poses accurately.
Figure 5: Registration of Teaser++ and ours. RRE/RTE metrics are shown, with RTE scaled by ×$\times 100$. The ground-truth Euler angles in this case are $\boldsymbol{\omega}{=}(202.97^{\circ},352.18^{\circ},256.89^{\circ})$.

### v2

@@ 3.2 Dataset Construction
As introduced in Eq. (1), we decompose the level-set parameters $\boldsymbol{\theta}$ into $\boldsymbol{\mu}+\Delta\boldsymbol{\theta}$. This is inspired by the reparameterization trick in variational autoencoders Kingma et al. (2019). It emulates a normal distribution with expectation $\boldsymbol{\mu}$ and homogeneous standard deviation 111. We follow this decomposition to construct a dataset of level-set parameters with shape transformations in two stages.
@@
with $\{\Delta\theta_{1}^{mn}\}$ being pose-dependenet. $\odot$ denotes the Hadamard product. The tanh function constrains $\Delta\boldsymbol{\theta}$ to be in the range [1,1][-1,1][ - 1 , 1 ]. We train the above parameters similarly based on Algorithm 1. The major difference from stage one is that all shapes in a batch become clones of the individual shape being fitted. In addition, note that $\boldsymbol{\theta}=\boldsymbol{\mu}+\Delta\boldsymbol{\theta}$ in the Algorithm in this stage.
@@
Why not meta-learning like MetaSDF? MetaSDF Sitzmann et al. (2020b) applies the meta-learning technique Finn et al. (2017) to per-category shape reconstruction of DeepSDF. It introduces a large number of additional parameters (i.e., per-parameter learning rates) to train the network, which is computation-intensive. Besides, its three gradient updates of the network during inference stage often result in unsatisfactory surface quality Chou et al. (2022). Moreover, the loss function we employ in Eq. (7) for unsupervised surface reconstruction requires input gradients of the SDF network to be computed with backpropagation at every iteration, prohibiting the application of MetaSDF or MAML Finn et al. (2017).
@@ 4.1 Encoder-based Semantic Learning
[…]
@@
[…]
The PECat block differentiate level-set parameters from different SDF layers in tensor $\Theta_{2}$. It applies positional encoding Mildenhall et al. (2020) to the layer index $\ell{\in}\{2,3,4,5,6,7\}$, and concatenates the encoded features $\text{PE}(\ell)$ to their corresponding layer parameters in $\Theta_{2}$. Let $d_{pe}$ be the dimension of $\text{PE}(\ell)$. Thus, the tensor $outA$ in Fig. 2(c) has dimension of $6{\times}256{\times}(257{+}d_{pe})$. We process it with BasePool2A. Besides, to extract semantics from the 2nd dimension of $\Theta_{2}$, we transpose $\Theta_{2}$ into $\Theta_{2}^{\intercal}{\in}{\mathbb{R}^{6{\times}257{\times}256}}$ and concatenate it with $\text{PE}(\ell)$ to get a tensor $outB$ of size $6{\times}257{\times}(256{+}d_{pe})$, processed by BasePool2B.
The BaseNet1 and BaseNet3 modules compute branch features $\mathbf{a}_{1}{\in}\mathbb{R}^{256}$ and $\mathbf{a}_{3}{\in}\mathbb{R}^{256}$, respectively, based on $\Theta_{1}$ and $\Theta_{3}$. Meanwhile, BasePool2A and BasePool2B each compute features of size 256256256, resulting in the concatenated feature $\mathbf{a}_{2}{\in}\mathbb{R}^{512}$. These features are further concatenated to form a unified global surface feature $\mathbf{s}_{g}{\in}\mathbb{R}^{1024}$, which is applicable to both shape classification and retrieval. We train the network with standard cross-entropy loss. The classifier comprises a single FC layer.
Our analysis considers level-set parameters from all layers of SDF, which forms a complete shape representation. In contrast, prior works utilized partial level-set parameters. For instance, Luigi et al. (2023) did not include the first layer parameters, while Erkoç et al. (2023) excluded the final layer parameters. Further, given our proposed decomposition of $\boldsymbol{\theta}$ and the fact that $\boldsymbol{\mu}$ is shared by all shapes, we normalize the level-set parameters with $\boldsymbol{\mu}$ and study shape semantics with $\Delta\boldsymbol{\theta}$, the instance parameters of each shape.
@@ 5 Experiment
ShapeNet Chang et al. (2015) and Manifold40 Hu et al. (2022) provide multi-category 3D shapes that are well-suited for geometric research in computer graphics and robotics. For the interested shape analysis, we construct Level-Set Parameter Data (LSPData) as continuous shape representations for these datasets. The two-stage training of HyperSE3-SDF in § 3.2 is adopted in this process. In the first stage of constructing the LSPData, we utilize 20 shapes from each class in ShapeNet and 7 shapes per class in Manifold40 to train the pose-dependent initialization $\boldsymbol{\mu}$. The point clouds of each shape for SDF are sampled from the shape surfaces in Manifold40 and sourced from previous work Mescheder et al. (2019) in ShapeNet. We did not consider the loudspeaker category in ShapeNet due to the intricate internal structures of the shapes.
@@ 5.1 Understand the Parameter Data
In the proposed dataset construction, we train HyperSE3-SDF for all shapes to create the LSPData with transformations for continuous shape analysis. To validate this approach, we conduct two extra experiments based on the plain SDF network to construct the LSPData, using different settings for $\boldsymbol{\mu}$. In the first experiment, we randomly initialize $\boldsymbol{\mu}$ and apply it to all shapes as Luigi et al. (2023). In the second, we learn $\boldsymbol{\mu}$ using the same training samples as those used for HyperSE3-SDF. The training process resembles the algorithm described in § 3.2, except that we fix the pose as $(\mathbf{I},\mathbf{0})$ without sampling.
In this ablation study, we compare the constructed LSPData for five major categories of ShapeNet Park et al. (2019a). We present the t-SNE embeddings Van der Maaten and Hinton (2008) of the data constructed by random and learned $\boldsymbol{\mu}$ in Fig. 3. We note that the embeddings for our HyperSE3-SDF in pose $(\mathbf{I},\mathbf{0})$ are similar to Fig. 3(b). It can be seen that the learned $\boldsymbol{\mu}$ significantly enhances shape semantics compared to the randomly initialized counterpart. Table 1 compares the shape classification results of the different LSPData under different rotation setups. We utilize the encoder-based network in § 4.1 for the classification. ‘$(\mathbf{I},\mathbf{0})$’ represents testing with data in reference poses. ‘$z$/SO(3)’ indicates training with data rotated around $z$ axis but testing on data arbitrarily rotated in SO(3) Esteves et al. (2018). We obtain the transformed LSPData of plain SDF by applying Eq. (4). Notably, the transformed LSPData from HyperSE3-SDF surpasses those from Euclidean computations, while the normalized LSPData $\Delta\boldsymbol{\theta}$ outperforms the unnormalized $\boldsymbol{\theta}$. We also conduct ablation studies to identify the optimal pose-dependent subset within the level-set parameters for the hypernetwork to generate. Details can be found in the supplementary material.
@@ 5.2 Shape Classification and Retrival
We apply the level-set parameters to semantic analysis on ShapeNet and Manifold40. For time concern, a maximum of 2000 shapes are reconstructed for each class in ShapeNet. We filter the continuous shapes based on their Chamfer distances to ensure surface quality. This results in a dataset of 17,656 shapes for ShapeNet and 10,859 for Manifold40. Figure 4 plots the mean and standard deviation of Chamfer distance (CD1) for each class in ShapeNet, illustrating the surface quality represented by our LSPData. Additionally, the mean and standard deviation of CD1 for all classes in Manifold40 are 0.71020.71020.7102 and 0.14430.14430.1443, respectively. Note that our chamfer distances are averaged across multiple random poses due to the introduced surface transformation. Besides, our training set in ShapeNet comprises 200 shapes per class, while in Manifold40, it contains 50% of the shapes per class, up to 200 shapes.
ShapeNet. We conduct pose-related semantic analysis on ShapeNet using rotation setups of $z$/SO(3) and $(\mathbf{I},\mathbf{0})$. Table 2 demonstrates the feasibility of shape analysis based on level-set parameters without dependence on point clouds and meshes Luigi et al. (2023). The proposed level-set parameters, LSPData, show comparable performance to point clouds in shape classification of pose $(\mathbf{I},\mathbf{0})$. In the challenging setup of $z$/SO(3), our encoder-based network for LSPData outperforms the rotation-equivariant network VN-DGCNN Deng et al. (2021) for point clouds. Conversely, the point cloud-based networks, PointNet Qi et al. (2017a) and DGCNN Wang et al. (2019a), yield unsatisfactory results for not addressing rotation equivariance. We note that the equivariance of VN-DGCNN comes with time cost, requiring 9×9$\times$9 × more time for training compared to DGCNN. In contrast, the classification network for LSPData converges rapidly (e.g., within 202020 minutes on the same machine using a single NVIDIA RTX 4090), akin to the observation in Dupont et al. (2022).
@@
Method pose OA plane bench cab car chair display lamp rifle sofa table phone vessel PointNet Qi et al. (2017a) $z$/SO(3) 27.34 17.01 33.16 27.41 13.87 29.83 21.88 35.28 61.11 12.36 11.74 22.61 38.13 DGCNN Wang et al. (2019a) 24.41 81.76 4.82 19.31 43.75 13.60 4.28 59.87 0.00 4.09 4.65 8.38 22.16 VN-DGCNN Deng et al. (2021) 90.40 97.01 81.84 93.35 98.35 86.98 83.62 89.05 96.17 89.39 80.95 92.95 93.13 LSPData ($\Delta\boldsymbol{\theta}$) 91.54 97.24 82.36 83.18 98.08 93.43 86.55 89.46 98.15 88.81 85.01 97.21 95.15 PointNet Qi et al. (2017a) ($\mathbf{I,0}$) 92.48 97.60 86.46 95.02 98.56 94.07 90.59 87.37 98.15 91.81 81.68 91.76 94.18 DGCNN Wang et al. (2019a) 94.55 99.00 84.35 95.64 99.86 93.94 94.25 93.31 99.15 92.06 88.26 97.87 97.69 LSPData ($\Delta\boldsymbol{\theta}$) 93.46 98.59 84.02 88.16 98.83 94.65 95.11 93.14 99.57 90.56 87.00 94.15 95.75
@@
Manifold40. We also compare the effectiveness of LSPData in shape classification and retrieval with point clouds under exhaustive rotation augmentations, i.e., the SO(3)/SO(3) setup, abbreviated as SO(3) in Table 3. For retrieval, we take features after the average pooling layer to match shapes in the feature space via Euclidean distances. We evaluate the retrieval performance using mean Average Precision (mAP) alongside the top-1/5/10 recalls. It can be noticed that the proposed encoder-based network for LSPData still outperforms the VN-DGCNN Deng et al. (2021), DGCNN Wang et al. (2019a), and PointNet Qi et al. (2017a) for point clouds. The performance gap between Manifold40 and ShapeNet is mainly attributed to the restricted number of shapes in certain classes of Manifold40.
@@ 5.3 Object Pose Estimation
Data preparation. To prepare the partial point clouds for pose estimation, we select three categories from ShapeNet: airplane, car, and chair, with 10 shapes utilized in each category. For every shape, we create 10 ground-truth transformations with random rotations in the range of $[0,2\pi]$ and translations in the range of [0.1,0.1][-0.1,0.1][ - 0.1 , 0.1 ]. For each transformed shape, we create the partial point cloud from its full point cloud representation, using hidden point removal Katz et al. (2007). This results in 300 pairs that covers partial and full point clouds with limited overlaps and large rotations. We introduce different levels of noise, $\sigma{\in}\{0.01,0.03\}$, and an outlier ratio of 30% to the partial point clouds for further challenges.
We compare the estimation quality of our method, Proposed V1 and V2 (see § 4.2), with optimization-based registration methods, including ICP Besl and McKay (1992), FGR Zhou et al. (2016), and TEASER++Yang et al. (2020) in Table 4. Their performance is evaluated based on the relative rotation/translation errors (RRE/RTE). See the supplementary for definitions of the two metrics. RRE is reported in degrees, and RTE is scaled by ×100$\times$ 100× 100. Notably, the proposed method effectively estimates poses with arbitrary rotations from partial-view point clouds, even in the presence of significant noise and outliers. Visualized examples are provided in Fig. 5, with more in the supplementary. We notice that while TEASER++ recovers most poses in the clean data, its performance drops with noise and outliers. ICP and FGR struggle with estimating large rotations even for clean point clouds. If a method fails in simpler settings, we cease testing it on more challenging data. The proposed method takes $\sim$50 seconds to estimate the poses accurately.
Figure 5: Registration of Teaser++ and ours. RRE/RTE metrics are shown, with RTE scaled by ×100$\times$ 100× 100. The ground-truth Euler angles in this case are $\boldsymbol{\omega}{=}(202.97^{\circ},352.18^{\circ},256.89^{\circ})$.